OpenStudy (anonymous):

Find the angle between the vectors u and v. u=cos(π/6)i+sin(π/6)j v=cos(3π/4)i+sin(3π/4)j A. 15° B. 75° C. 95° D. 105° E. 135°

OpenStudy (anonymous):

D

OpenStudy (anonymous):

Ok I know it's alot to write out but could explain how you found it because I found u.v but after that I got lost... I know the equation is: \[u.v/\left| u \right|\left| v \right|\]

OpenStudy (anonymous):

FIND THE DOT PRODUCT WHICH WILL GIVE U "COSINE FORMULA. COSA COSB + SINA SINB= COS(A-B)

OpenStudy (anonymous):

ok thanks